# BEE-RAG: Balanced Entropy Engineering for Retrieval-Augmented Generation

### v1

@@ Introduction
Retrieval-augmented generation (RAG) has emerged as a transformative paradigm for augmenting large language models (LLMs) with dynamically integrated external knowledge (Gao et al. 2023; Lewis et al. 2020). While conventional RAG systems demonstrate remarkable capabilities in knowledge-intensive tasks, their performance exhibits critical fragility when processing extended contextual inputs (Ren et al. 2023). This vulnerability arises from the inherent requirement of RAG frameworks to incorporate external retrieved documents, which typically results in substantially longer input sequences compared to standard generation tasks (Zhang et al. 2024a). This limitation manifests acutely in scenarios demanding long-context comprehension, such as scientific literature analysis, multi-hop reasoning, and cross-document synthesis, where the growing complexity of modern knowledge systems necessitates robust processing of extended text spans (Liu et al. 2023).
@@
Typically, existing efforts to mitigate this challenge focus predominantly on heuristic document filtering, extrapolative training, or architectural modifications, introducing worth noting trade-offs. Threshold-based retrieval document truncation or selection reduces context length at the cost of discarding potentially relevant information (Jeong et al. 2024; Wang et al. 2024). While extended context training may enhance RAG performance (Lin et al. 2024; Zhang et al. 2024a), it substantially increases computational demands and risks compromising model generalization. The introduction of auxiliary modules, such as specialized encoders (Yen, Gao, and Chen 2024), further escalates the complexity of the system.
@@
Task Formulation.  The task of retrieval-augmented generation (RAG) involves the integration of retrieval-based techniques with generative models to enhance the accuracy and coherence of text generation tasks (Lewis et al. 2020). A typical RAG system consists of a corpus $C$, a retriever $R$, and a generative model $LLM$. Given a user query or input context $q$, the retriever $R$ retrieves relevant documents $D_{\text{q}}$ from the corpus. The generative model $LLM$ uses both the query $q$ and the retrieved documents to generate a response:
@@
Entropy Engineering.  In this work, we reconsider RAG from an information-theoretic perspective and introduce the concept of Entropy Engineering. In the standard Transformer (Waswani et al. 2017), the attention is computed as:
@@
Based on this, inspired by existing analyses (Zhang et al. 2024c), we measure the amount of information a position receives using discrete entropy, as shown in the following equation:
@@
We assume that each row of the query matrix $\mathbf{Q}\in\mathbb{R}^{n\times d}$ and key matrix $\mathbf{K}\in\mathbb{R}^{n\times d}$ is an independent, isotropic sub-Gaussian vector with zero mean and identity covariance, i.e., $\mathbb{E}[\mathbf{q}_{i}]=\mathbb{E}[\mathbf{k}_{i}]=\mathbf{0}$. By the high-dimensional central limit theorem (Bolthausen and Wüthrich 2013), the scaled dot product $\xi_{i,j}=\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d}}$ converges in distribution to a standard Gaussian, i.e., $\xi\xrightarrow{d}\mathcal{N}(0,1)$. We further assume that the document-level importance bias follows a Gaussian distribution $\beta\sim\mathcal{N}(\mu,\sigma^{2})$ and is independent of $\xi$. This independence reflects the fact that the same token may appear in documents of varying importance, and thus the semantic similarity (captured by $\xi$) and document relevance (captured by $\beta$) should be modeled as separate random variables. Thus, we have:
@@
Figure 2: Based on Equation 9, this illustrates how to set $\mu$ and $\sigma$ to balance the entropy increase caused by growing context length $n$.
The solution to Equation (9) is shown in Figure 2. When $\mu$ is set to a small value, $\sigma$ approximates a linear relationship with $n$. Therefore, we need to constrain $\sigma$ to increase as $n$ grows, ensuring that the context entropy remains stable as the context length increases. Our experimental results confirm that the parameters derived from theory are consistent with those selected empirically.
@@
Inspired by the success of prompt engineering for LLMs in retrieval and ranking tasks (Zhuang et al. 2024; Pradeep, Sharifymoghaddam, and Lin 2023), we use a prompt-induced attention calibration strategy to obtain importance scores for each document. Specifically, we append a prompt to each document for obtaining hidden state representations of importance. To extract importance scores from these representations, we apply an early-decoding strategy (Cammarata et al. 2020), using lm_head(·) to map the generation probability of a critic token as the document’s importance score at each layer. For instance, the prompt can be “Does the passage support the answer to the question?”, the critic token is “yes”).
@@
The cross-attention mechanisms employed in LLMs may inadvertently influence importance score estimation for individual documents, as it enables inter-document interference through shared attention parameters. Inspired by previous work (Ratner et al. 2023), we develop a parallel scoring mechanism in RAG architecture through attention mask modifications. This enables simultaneous computation of document importance and task outputs in a single forward pass. In this way, we create independent processing streams that prevent cross-document contamination, which is a critical foundation for stabilizing the context entropy.
@@
The core of our approach lies in the specially designed efficient vector fine-tuning method, which bridges token-level and sentence-level representations. Specifically, inspired by previous work (Wu et al. 2025), we transforms token-level vectors into sentence-level vectors:
@@
During implementation, we observe unconstrained values of $\bm{\beta}$ leading to gradient explosion. To address it, we employ orthogonal initialization (Saxe, McClelland, and Ganguli 2013) on $\bm{R}$ and implement additional scaling on $\bm{\beta}$ when its values exceed a certain threshold.
@@ Discussion
Novelty and Effectiveness.  Our key innovation lies in introducing an additive balancing entropy factor into the attention mechanism, achieving two critical objectives: (1) maintaining stable context entropy across varying input lengths through mean and variance constraints, thereby enhancing performance in long-context scenarios; and (2) adaptively guiding the LLM to focus on crucial segments, improving its ability to effectively utilize knowledge. Unlike existing attention reweighting approaches that either uniformly reduce context entropy (Su 2021) (hindering important information extraction) or ignore attention entropy (Yan et al. 2024) (leading to unstable performance across varying context lengths). We further introduce novel adaptations for zero-shot and parameter-efficient fine-tuning scenarios: a prompt-induced importance scoring mechanism with parallel scoring-generation architecture, and an efficient vector fine-tuning strategy. These components collectively enhance BEE-RAG’s accuracy in question-answering tasks
Efficiency.  We further analyze the efficiency of our EE-RAG framework. The parallel mechanism handles $\beta$ calculation and answer generation simultaneously. Due to the quadratic time complexity in the transformer mechanism, our design effectively controls computational overhead, achieving performance comparable to vanilla RAG with improved scalability for longer context. These efficiency gains result from our optimized attention complexity handling.
@@
We conduct experiments on both single-hop and multi-hop question-answering (QA) tasks. Specifically, we evaluate our approach on the Natural Questions (NQ) (Kwiatkowski et al. 2019), TriviaQA (Joshi et al. 2017), HotpotQA (Yang et al. 2018), and 2WikiMultihopQA (2WikiQA) (Ho et al. 2020) datasets.
@@
We use Exact Match (EM(Lee, Chang, and Toutanova 2019) and F1 (Karpukhin et al. 2020) to assess the QA accuracy of retrieval-augmented LLMs, which are commonly adopted in QA evaluation. EM checks if predicted answers exactly match the ground truth, while F1 measures their overlap in precision and recall.
@@
We select LLaMA-3-8B (Dubey et al. 2024) and Qwen-2.5-7B (Yang et al. 2024) as base LLM. For the zero-shot setting (Section Intrinsic Multi-Importance Inference), we derive the theoretical values for $\mu$ and $\sigma^{2}$ based on Equation 9, and perform a search around these values to determine their optimal configuration. For the adaptive balancing entropy factor learning, we set $d_{r}=8$. We conduct the experiments on 8 NVIDIA A100 40GB GPUs, with a learning rate of 5e-4, a cosine learning rate scheduler, and a batch size of 16. To demonstrate the robustness of our training approach, we train the LLMs for no more than 500 steps solely on the NQ dataset and subsequently validate its performance across multiple datasets.
@@
• Zero-shot methods. We select several prompt-based strategies that enhance generation performance in RAG scenarios for comparison, including prompt-based methods (Chain-of-Thought (Wei et al. 2022), Chain-of-Note (Yu et al. 2023), Self-Critic (Asai et al. 2023)), architecture-driven methods (PCW (Ratner et al. 2023), Position-Engineering (He et al. 2024)), and attention reweighting method (Multiply-Attention (Chiang and Cholak 2022)).
• Parametric-efficient fine-tuning methods: We choose two commonly used PEFT approaches in RAG, including LoRA (Hu et al. 2022) and Prefix-Tuning (Li and Liang 2021) for comparison with our parameter-efficient training strategy.
@@
Table 1 presents the results of our approach and the baselines on four RAG tasks of open-domain QA. It can be observed that:
@@
In this section, we further present an analysis from two aspects: the scaling of LLMs and the quality of the retrieved documents.
@@
Recent efforts have aimed to improve RAG performance under long contexts (Laban et al. 2024). PCW (Ratner et al. 2023) restricts attention within fixed-size windows to reduce attention dilution. CEPE (Yen, Gao, and Chen 2024) uses a lightweight encoder to process inputs chunk by chunk, enabling better use of long contexts via cross-attention. RAFT (Zhang et al. 2024b) and Chain-of-Notes (Yu et al. 2023) extend the context capacity of models through instruction tuning on long-input data. However, these methods either yield limited performance gains or incur high training costs. Some works explore the effect of attention distribution (Wang et al. 2023). Su (2021) adjust attention scaling by context length to match short-context behavior, but this fails to highlight key segments and may miss important information.

### v2

@@ Introduction
Retrieval-augmented generation (RAG) has emerged as a transformative paradigm for augmenting large language models (LLMs) with dynamically integrated external knowledge (wang2025unveiling; lewis2020retrieval). While conventional RAG systems demonstrate remarkable capabilities in knowledge-intensive tasks, their performance exhibits critical fragility when processing extended contextual inputs (ren2023investigating). This vulnerability arises from the inherent requirement of RAG frameworks to incorporate external retrieved documents, which typically results in substantially longer input sequences compared to standard generation tasks (zhang2024soaring; li2025analyzing). This limitation manifests acutely in scenarios demanding long-context comprehension, such as scientific literature analysis, multi-hop reasoning, and cross-document synthesis, where the growing complexity of modern knowledge systems necessitates robust processing of extended text spans (liu2023lost).
@@
Typically, existing efforts to mitigate this challenge focus predominantly on heuristic document filtering, extrapolative training, or architectural modifications, introducing worth noting trade-offs. Threshold-based retrieval document truncation or selection reduces context length at the cost of discarding potentially relevant information (jeong2024adaptive; wang-etal-2024-rear). While extended context training may enhance RAG performance (linra; tang2025enhancing), it substantially increases computational demands and risks compromising model generalization (zhan2024over). The introduction of auxiliary modules, such as specialized encoders (DBLP:conf/acl/YenG024), further escalates the complexity of the system (ren2025llm).
@@
Task Formulation.  The task of retrieval-augmented generation (RAG) involves the integration of retrieval-based techniques with generative models to enhance the accuracy and coherence of text generation tasks (lewis2020retrieval). A typical RAG system consists of a corpus $C$, a retriever $R$, and a generative model $LLM$. Given a user query or input context $q$, the retriever $R$ retrieves relevant documents $D_{\text{q}}$ from the corpus. The generative model $LLM$ uses both the query $q$ and the retrieved documents to generate a response:
@@
Entropy Engineering.  In this work, we reconsider RAG from an information-theoretic perspective and introduce the concept of Entropy Engineering. In the standard Transformer (waswani2017attention), the attention is computed as:
@@
Based on this, inspired by existing analyses (zhang2024attention), we measure the amount of information a position receives using discrete entropy, as shown in the following equation:
@@
We assume that each row of the query matrix $\mathbf{Q}\in\mathbb{R}^{n\times d}$ and key matrix $\mathbf{K}\in\mathbb{R}^{n\times d}$ is an independent, isotropic sub-Gaussian vector with zero mean and identity covariance, i.e., $\mathbb{E}[\mathbf{q}_{i}]=\mathbb{E}[\mathbf{k}_{i}]=\mathbf{0}$. By the high-dimensional central limit theorem (bolthausen2013bernoulli), the scaled dot product $\xi_{i,j}=\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d}}$ converges in distribution to a standard Gaussian, i.e., $\xi\xrightarrow{d}\mathcal{N}(0,1)$. We further assume that the document-level importance bias follows a Gaussian distribution $\beta\sim\mathcal{N}(\mu,\sigma^{2})$ and is independent of $\xi$. This independence reflects the fact that the same token may appear in documents of varying importance, and thus the semantic similarity (captured by $\xi$) and document relevance (captured by $\beta$) should be modeled as separate random variables. Thus, we have:
@@
Based on the solution to Equation (9), we need to constrain $\sigma$ to increase as $n$ grows, ensuring that the context entropy remains stable as the context length increases. Our experimental results confirm that the parameters derived from theory are consistent with those selected empirically.
@@
Inspired by the success of prompt engineering for LLMs in retrieval and ranking tasks (zhuang2024setwise; pradeep2023rankvicuna), we use a prompt-induced attention calibration strategy to obtain importance scores for each document. Specifically, we append a prompt to each document for obtaining hidden state representations of importance. To extract importance scores from these representations, we apply an early-decoding strategy (cammarata2020thread:), using lm_head(·) to map the generation probability of a critic token as the document’s importance score at each layer. For instance, the prompt can be “Does the passage support the answer to the question?”, the critic token is “yes”).
@@
The cross-attention mechanisms employed in LLMs may inadvertently influence importance score estimation for individual documents, as it enables inter-document interference through shared attention parameters. Inspired by previous work (ratner2023parallel), we develop a parallel scoring mechanism in RAG architecture through attention mask modifications. This enables simultaneous computation of document importance and task outputs in a single forward pass. In this way, we create independent processing streams that prevent cross-document contamination, which is a critical foundation for stabilizing the context entropy.
@@
The core of our approach lies in the specially designed efficient vector fine-tuning method, which bridges token-level and sentence-level representations. Specifically, inspired by previous work (wu2025reft), we transforms token-level vectors into sentence-level vectors:
@@
During implementation, we observe unconstrained values of $\bm{\beta}$ leading to gradient explosion. To address it, we employ orthogonal initialization (saxe2013exact) on $\bm{R}$ and implement additional scaling on $\bm{\beta}$ when its values exceed a certain threshold.
@@ Discussion
Novelty and Effectiveness.  Our key innovation lies in introducing an additive balancing entropy factor into the attention mechanism, achieving two critical objectives: (1) maintaining stable context entropy across varying input lengths through mean and variance constraints, thereby enhancing performance in long-context scenarios; and (2) adaptively guiding the LLM to focus on crucial segments, improving its ability to effectively utilize knowledge. Unlike existing attention reweighting approaches that either uniformly reduce context entropy (kexuefm-8823) (hindering important information extraction) or ignore attention entropy (yan2024corrective) (leading to unstable performance across varying context lengths). We further introduce novel adaptations for zero-shot and parameter-efficient fine-tuning scenarios: a prompt-induced importance scoring mechanism with parallel scoring-generation architecture, and an efficient vector fine-tuning strategy. These components collectively enhance BEE-RAG’s accuracy in question-answering tasks
Efficiency.  We further analyze the efficiency of our EE-RAG framework. The parallel mechanism handles $\beta$ calculation and answer generation simultaneously. Due to the quadratic time complexity in the transformer mechanism, our design effectively controls computational overhead, achieving performance comparable.
@@
We conduct experiments on both single-hop and multi-hop question-answering (QA) tasks. Specifically, we evaluate our approach on the Natural Questions (NQ) (kwiatkowski2019natural), TriviaQA (joshi2017triviaqa), HotpotQA (Yang2018HotpotQAAD), and 2WikiMultihopQA (2WikiQA) (ho2020constructing) datasets.
@@
We use Exact Match (EM(lee2019latent) and F1 (dpr2020) to assess the QA accuracy of retrieval-augmented LLMs, which are commonly adopted in QA evaluation. EM checks if predicted answers exactly match the ground truth, while F1 measures their overlap in precision and recall.
@@
We select LLaMA-3-8B (dubey2024llama) and Qwen-2.5-7B (yang2024qwen2) as base LLM. For the zero-shot setting (Section Intrinsic Multi-Importance Inference), we derive the theoretical values for $\mu$ and $\sigma^{2}$ based on Equation 9, and perform a search around these values to determine their optimal configuration. For the adaptive balancing entropy factor learning, we set $d_{r}=8$. We conduct the experiments on 8 NVIDIA A100 40GB GPUs, with a learning rate of 5e-4, a cosine learning rate scheduler, and a batch size of 16. To demonstrate the robustness of our training approach, we train the LLMs for no more than 500 steps solely on the NQ dataset and subsequently validate its performance across multiple datasets.
@@
• Zero-shot methods. We select several prompt-based strategies that enhance generation performance in RAG scenarios for comparison, including prompt-based methods (Chain-of-Thought (wei2022chain), Chain-of-Note (yu2023chain), Self-Critic (asaiself)), architecture-driven methods (PCW (ratner2023parallel), Position-Engineering (he2024position)), and attention reweighting method (Multiply-Attention (chiang2022overcoming)).
• Parametric-efficient fine-tuning methods: We choose two commonly used PEFT approaches in RAG, including LoRA (hulora) and Prefix-Tuning (li2021prefix) for comparison with our parameter-efficient training strategy.
@@
Table 1 presents the results of our approach and the baselines on four RAG tasks of open-domain QA.
@@
In this section, we further present an analysis from two aspects: the scaling of LLMs and the quality of the documents.
@@
Recent efforts have aimed to improve RAG performance under long contexts (laban2024summary; wang2025reinforced). PCW (ratner2023parallel) restricts attention within fixed-size windows to reduce attention dilution. CEPE (DBLP:conf/acl/YenG024) uses a lightweight encoder to process inputs chunk by chunk, enabling better use of long contexts via cross-attention. RAFT (zhang2024raft) and Chain-of-Notes (yu2023chain) extend the context capacity of models through instruction tuning on long-input data. However, these methods either yield limited performance gains or incur high training costs. Some works explore the effect of attention distribution (wang2023label). kexuefm-8823 adjust attention scaling by context length to match short-context behavior, but this fails to highlight key segments.